What's the absolute value of -144
Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

Ans: 144
---
Absolute values are never negative.
|a-b| means "the distance from a to b.
|a-b| = |b-a|
-----------------
|-144| = |-144-0| = 144
===========================
Cheers,
Stan H.
=========

Answer by ewatrrr(24785)


Hi
|-144| = 144 (the numerical value of the number)